NERSCPowering Scientific Discovery Since 1974

March 15

Attending:  Jay and Eric from NERSC;  users Craig and Andrei

Cluster Status:  Still running at full capacity

Recent Outages:  SWO Sunday.  Problem was corruption on /common which was restricted to dayabay files.  Did rsync of ~25% of /home also.

Upcoming downtimes:  Nothing for PDSF but NGF downtime 3/30-3/31.  Dayabay discussed their need for webhosting and how this will impact their work.

Procurements:  Small procurement of local disks for STAR;  dayabay might want 20 or 40 TB in the near future.

Interactive Node abuse:  Memory limits are in place;  users running too many jobs, too long jobs; we don't want to auto-kill as it is more difficult with NX being used a lot

Grid issues:  ATLAS needs permissions changed;  server doesn't seem to be working right...

NXproxy:  Yushu is looking into how to provide this or an equivalent